# Merge-of-Thought Distillation

### v1

@@ 1 Introduction
As large language models (LLMs) with long chain-of-thought (CoT) capabilities continue to emerge Jaech et al. (2024); Yang et al. (2025a); Guo et al. (2025), reasoning distillation is becoming the key pathway for converting expensive reasoning ability into deployable efficiency. Compared with imitating only final answers, directly supervising the reasoning trajectory enables a smaller student model to learn multi-step solution procedures Luo et al. (2025b); Qin et al. (2025); Guo et al. (2025).
Building on this, the research focus is naturally shifting from “stacking quantity” to “improving quality and supporting mechanisms”: on the one hand, Muennighoff et al. (2025) performs SFT on only 1,000 teacher-distilled samples and shows reasoning gains with test-time compute; on the other hand, Ye et al. (2025) shows that when pretraining already embeds rich mathematical knowledge, a few hundred carefully curated examples can effectively elicit complex reasoning. Overall, these advances jointly emphasize that efficient distillation of long CoT reasoning is a scalable route to high-accuracy small models.
However, real-world production rarely feature a “single oracle teacher.” We often have multiple candidate teacher LLMs and a growing pool of distilled CoT data, giving rise to a basic question: given a student model, how should we pick the most suitable teacher? Empirically, teacher choice matters—the teacher can imprint a recognizable “style signature” on the student Chen et al. (2025b); mismatches between teacher and student can weaken the transfer of long CoT skills Wu et al. (2025b). As illustrated in Figure 1, our observations are consistent: different students have different “best teachers,” and even for the same student the best teacher can vary across datasets. Such phenomena challenge the naive assumption that “a bigger/stronger teacher is necessarily better,” prompting us to consider: rather than centering on a single teacher and taking cost on teacher selection, maybe we should transfer across multiple teachers.
@@
A natural follow-up question is: can we fuse the strengths of multiple teachers—integrating diverse reasoning styles and complementary skills into a single student to achieve stronger performance? Recent studies have shown that long CoTs, due to excessive length and redundant steps, tend to accumulate more noise and irrelevant content Luo et al. (2025a); Zhang et al. (2025); Li et al. (2025b). It is unclear whether, in mixed-teacher long-CoT distillation, such noise is amplified through interactions, and how to suppress idiosyncratic noise while preserving the consensus structure. These trends suggest that diversity of teachers and reasoning paths is an asset—provided we can overcome conflicts among various teachers’ supervision.
Model merging has already demonstrated that merging can help models fit to different data distributions and is widely applied to joint training across diverse domains and tasks Yu et al. (2024b); Zhou et al. (2024); Yadav et al. (2024). Inspired by these advantages, we aim to leverage model merging to overcome conflicts among various teachers’ supervision and, through continuous merge-and-training iterations, unify different teachers’ reasoning abilities and ultimately converge to a consensus reasoning landscape.
@@ 2 Related Work
Research on distilling long chains of thought (CoT) has progressed rapidly Wu et al. (2025b); Guo et al. (2025). Early work Li et al. (2023) showed that even small models can benefit from teacher CoT prompting and highlighted the importance of using diverse teachers to obtain varied reasoning chains. Skip-Thinking partitions long reasoning into coherent segments and skips non-essential parts, improving efficiency and alleviating gradient over-smoothing Chen et al. (2025a). Subsequent approaches like DLCoT Luo et al. (2025b) and KPOD Feng et al. (2024) further segment and simplify CoTs, employ keypoint weighting, and use progressive distillation to focus on critical tokens and steps. Studies on the key factors of CoT distillation reveal that teacher diversity and rationale granularity often have a greater impact than raw teacher accuracy Chen et al. (2025b). Recent works show that long-CoT capability can be bootstrapped with a handful of in-context examples Pang et al. (2025), distilled as summaries to improve long-context memory Ma et al. (2025), or integrated with vision reasoning using agent-based approaches Shi et al. (2024). These findings underscore that long-CoT distillation not only requires carefully curated examples but also faces challenges such as teacher selection, noise amplification and distillation efficiency. Nevertheless, most existing methods focus on a single teacher distillation; our work instead extends this line of work by fusing multiple teachers’ reasoning abilities into a single student to achieve stronger performance.
@@
Model merging fuses the parameters of multiple trained models into a single model, which is distinct from output-level ensembles Yang et al. (2024); Tam et al. (2024). Empirical studies show that merging tends to balance performance and safety better than mixing data across tasks or languages Yang et al. (2025b); Yadav et al. (2024); Yu et al. (2024b); Jin et al. . More advanced techniques adapt merging to pre-trained models by disentangling weights into magnitude and direction Yu et al. (2024a). Other approaches merge checkpoints during pre-training for faster convergence or use activation importance to retain critical parameters Li et al. (2025a); Nobari et al. (2025). Model merging has also been applied to combine models with different reasoning strategies and to merge heterogeneous architectures Wu et al. (2025a); Zhang et al. (2024). However, most existing work focuses on merging models specialised for different domains and tasks; by contrast, our approach merges student models distilled by different teachers on the same dataset to unify their reasoning abilities without conflicts among different teachers.
@@ 4.1 Datasets
We work in a one-question–multiple-answers (1Q–multiA) setting. We use two high-quality open-source mathematical datasets (BOBA inclusionAI (2025) and S1K Muennighoff et al. (2025) as our source datasets. From each source dataset, we sample 200 prompts and denote the resulting subsets as BOBA-200 and S1K-200. For every prompt, we query four teacher models—QWEN3-32B Yang et al. (2025a), QWQ Team (2024), DEEPSEEK-R1 Guo et al. (2025), and QWEN3-235B Yang et al. (2025a). Each teacher generates 16 responses with temperature set to 0.6 and max_tokens set to 32,768. For distillation, we randomly select one correct reasoning path among the 16 as the training label; if none of the 16 responses is correct, we discard that prompt for the corresponding teacher’s distillation corpus.
@@ 4.2 Training Configuration
We fine-tune QWEN3-8B, QWEN3-14B, and QWEN3-30-A3B Yang et al. (2025a) as base models across all experiments. We use a initial learning rate of 1e-5 and a global training batch size of 64. We evaluate the capabilities of the model in mathematical reasoning using AIME24 Math-AI (2024) and AIME25 Math-AI (2025).
@@ 5.2 MOT mitigates catastrophic forgetting and strengthens general reasoning
To assess whether CoT-style training with MOT affects basic capabilities, we evaluate the final checkpoints trained by MOT and by STD with the per-setting best teacher (Best STD) against the Base models on nine benchmarks: CEVAL (CEV) Seifert et al. (2024), SUPER_GPQA (SG) Du et al. (2025), SIMPLE_QA (SQ) Wei et al. (2024), IFEVAL (IFE) Zhou et al. (2023), MMLU_PRO (MP) Wang et al. (2024), MMLU_REDUX (MR) Gema et al. (2025), PhyBench (PB) Meng et al. (2024), LiveCodeBench (LCB) Jain et al. (2024), and GPQA-Diamond (GPQA-D) Rein et al. (2024). We group these benchmarks into three categories:

### v2

@@ 1 Introduction
As large language models (LLMs) with long chain-of-thought (CoT) capabilities continue to emerge (Jaech et al., 2024; Yang et al., 2025a; Guo et al., 2025), reasoning distillation is becoming the key pathway for converting expensive reasoning ability into deployable efficiency. Compared with imitating only final answers, directly supervising the reasoning trajectory enables a smaller student model to learn multi-step solution procedures (Luo et al., 2025b; Qin et al., 2025; Guo et al., 2025).
Building on this, the research focus is naturally shifting from “stacking quantity” to “improving quality and supporting mechanisms”: on the one hand, (Muennighoff et al., 2025) performs SFT on only 1,000 teacher-distilled samples and shows reasoning gains with test-time compute; on the other hand, (Ye et al., 2025) shows that when pretraining already embeds rich mathematical knowledge, a few hundred carefully curated examples can effectively elicit complex reasoning. Overall, these advances jointly emphasize that efficient distillation of long CoT reasoning is a scalable route to high-accuracy small models.
However, real-world production rarely feature a “single oracle teacher.” We often have multiple candidate teacher LLMs and a growing pool of distilled CoT data, giving rise to a basic question: given a student model, how should we pick the most suitable teacher? Empirically, teacher choice matters—the teacher can imprint a recognizable “style signature” on the student (Chen et al., 2025b); mismatches between teacher and student can weaken the transfer of long CoT skills (Wu et al., 2025b). As illustrated in Figure 1, our observations are consistent: different students have different “best teachers,” and even for the same student the best teacher can vary across datasets. Such phenomena challenge the naive assumption that “a bigger/stronger teacher is necessarily better,” prompting us to consider: rather than centering on a single teacher and taking cost on teacher selection, maybe we should transfer across multiple teachers.
@@
A natural follow-up question is: can we fuse the strengths of multiple teachers—integrating diverse reasoning styles and complementary skills into a single student to achieve stronger performance? Recent studies have shown that long CoTs, due to excessive length and redundant steps, tend to accumulate more noise and irrelevant content (Luo et al., 2025a; Zhang et al., 2025; Li et al., 2025b). It is unclear whether, in mixed-teacher long-CoT distillation, such noise is amplified through interactions, and how to suppress idiosyncratic noise while preserving the consensus structure. These trends suggest that diversity of teachers and reasoning paths is an asset—provided we can overcome conflicts among various teachers’ supervision.
Model merging has already demonstrated that merging can help models fit to different data distributions and is widely applied to joint training across diverse domains and tasks (Yu et al., 2024b; Zhou et al., 2024; Yadav et al., 2024). Inspired by these advantages, we aim to leverage model merging to overcome conflicts among various teachers’ supervision and, through continuous merge-and-training iterations, unify different teachers’ reasoning abilities and ultimately converge to a consensus reasoning landscape.
@@ 2 Related Work
Research on distilling long chains of thought (CoT) has progressed rapidly (Wu et al., 2025b; Guo et al., 2025). Early work (Li et al., 2023) showed that even small models can benefit from teacher CoT prompting and highlighted the importance of using diverse teachers to obtain varied reasoning chains. Skip-Thinking partitions long reasoning into coherent segments and skips non-essential parts, improving efficiency and alleviating gradient over-smoothing (Chen et al., 2025a). Subsequent approaches like DLCoT (Luo et al., 2025b) and KPOD (Feng et al., 2024) further segment and simplify CoTs, employ keypoint weighting, and use progressive distillation to focus on critical tokens and steps. Studies on the key factors of CoT distillation reveal that teacher diversity and rationale granularity often have a greater impact than raw teacher accuracy (Chen et al., 2025b). Recent works show that long-CoT capability can be bootstrapped with a handful of in-context examples (Pang et al., 2025), distilled as summaries to improve long-context memory (Ma et al., 2025), or integrated with vision reasoning using agent-based approaches (Shi et al., 2024). These findings underscore that long-CoT distillation not only requires carefully curated examples but also faces challenges such as teacher selection, noise amplification and distillation efficiency. Nevertheless, most existing methods focus on a single teacher distillation; our work instead extends this line of work by fusing multiple teachers’ reasoning abilities into a single student to achieve stronger performance.
@@
Model merging fuses the parameters of multiple trained models into a single model, which is distinct from output-level ensembles (Yang et al., 2024; Tam et al., 2024). Empirical studies show that merging tends to balance performance and safety better than mixing data across tasks or languages (Yang et al., 2025b; Yadav et al., 2024; Yu et al., 2024b; Jin et al., ). More advanced techniques adapt merging to pre-trained models by disentangling weights into magnitude and direction (Yu et al., 2024a). Other approaches merge checkpoints during pre-training for faster convergence or use activation importance to retain critical parameters (Li et al., 2025a; Nobari et al., 2025). Model merging has also been applied to combine models with different reasoning strategies and to merge heterogeneous architectures (Wu et al., 2025a; Zhang et al., 2024). However, most existing work focuses on merging models specialised for different domains and tasks; by contrast, our approach merges student models distilled by different teachers on the same dataset to unify their reasoning abilities without conflicts among different teachers.
@@ 4.1 Datasets
We work in a one-question–multiple-answers (1Q–multiA) setting. We use two high-quality open-source mathematical datasets (BOBA (inclusionAI, 2025) and S1K (Muennighoff et al., 2025) as our source datasets. From each source dataset, we sample 200 prompts and denote the resulting subsets as BOBA-200 and S1K-200. For every prompt, we query four teacher models—QWEN3-32B (Yang et al., 2025a), QWQ (Team, 2024), DEEPSEEK-R1 (Guo et al., 2025), and QWEN3-235B (Yang et al., 2025a). Each teacher generates 16 responses with temperature set to 0.6 and max_tokens set to 32,768. For distillation, we randomly select one correct reasoning path among the 16 as the training label; if none of the 16 responses is correct, we discard that prompt for the corresponding teacher’s distillation corpus.
@@ 4.2 Training Configuration
We fine-tune QWEN3-8B, QWEN3-14B, and QWEN3-30-A3B (Yang et al., 2025a) as base models across all experiments. We use a initial learning rate of 1e-5 and a global training batch size of 64. We evaluate the capabilities of the model in mathematical reasoning using AIME24 (Math-AI, 2024) and AIME25 (Math-AI, 2025).
@@ 5.2 MOT mitigates catastrophic forgetting and strengthens general reasoning
To assess whether CoT-style training with MOT affects basic capabilities, we evaluate the final checkpoints trained by MOT and by STD with the per-setting best teacher (Best STD) against the Base models on nine benchmarks: CEVAL (CEV) (Seifert et al., 2024), SUPER_GPQA (SG) (Du et al., 2025), SIMPLE_QA (SQ) (Wei et al., 2024), IFEVAL (IFE) (Zhou et al., 2023), MMLU_PRO (MP) (Wang et al., 2024), MMLU_REDUX (MR) (Gema et al., 2025), PhyBench (PB) (Meng et al., 2024), LiveCodeBench (LCB) (Jain et al., 2024), and GPQA-Diamond (GPQA-D) (Rein et al., 2024). We group these benchmarks into three categories: